# SeBS-Flow: Benchmarking Serverless Cloud Function Workflows

### v2

@@ VIII Related Work
Wen et al. [74] conducts a performance investigation of serverless workflows using two applications and microbenchmarks with varying numbers of functions, payload size, and parallelism. While they measure the execution time and estimate overhead, they do not evaluate scalability, billing, or investigate overhead sources. Instead, we focus on a wider collection of applications and propose a unifying model that allows developers to deploy and evaluate a single implementation across many cloud platforms. Moreover, we make all benchmark codes available and provide a ready-to-use benchmarking platform. Finally, we evaluated serverless Google Cloud Workflows instead of the non-serverless Google Cloud Composer.
@@
Other authors analyzed the productivity of workflow languages and proposed alternative models. AFCL [75] is a custom and provider-independent orchestration language for serverless workflows, implemented on top of AWS Step Functions and IBM Composer. Burckhardt et al. explore the semantics of Durable Functions [23] and propose Netherite [76], a new engine to replace Azure Durable Functions.
IX Conclusions
@@ Acknowledgments
Larissa Schmid is supported by the pilot program Core Informatics of the Helmholtz Association (HGF). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 program (grant agreement PSAP, No. 101002047). We would also like to thank the Swiss National Supercomputing Centre (CSCS) for providing us with access to their HPC machine Ault. We thank Amazon Web Services for supporting this research with credits through the AWS Cloud Credit for Research, and Google Cloud Platform through the Google Cloud Research Credits program with the award GCP19980904.

### v3

@@ 8. Related Work
Wen et al. (Wen and Liu, 2021) conducts a performance investigation of serverless workflows using two applications and microbenchmarks with varying numbers of functions, payload size, and parallelism. While they measure the execution time and estimate overhead, they do not evaluate scalability, billing, or investigate overhead sources. Instead, we focus on a wider collection of applications and propose a unifying model that allows developers to deploy and evaluate a single implementation across many cloud platforms. Moreover, we make all benchmark codes available and provide a ready-to-use benchmarking platform. Finally, we evaluated serverless Google Cloud Workflows instead of the non-serverless Google Cloud Composer. XFBench (Kulkarni et al., 2024) provides chaining of different functions and deploying them to AWS Step Functions and Azure Durable Functions, while we focus on realistic and complete applications. Moreover, they do not consider cloud-native data movement between functions via cloud storage, do not evaluate the overhead of their platform transcription, and can not compare pricing between platforms.
@@
Other authors analyzed the productivity of workflow languages and proposed alternative models. AFCL (Ristov et al., 2021) is a custom and provider-independent orchestration language for serverless workflows, implemented on top of AWS Step Functions and IBM Composer. Burckhardt et al. explore the semantics of Durable Functions (Burckhardt et al., 2021) and propose Netherite (Burckhardt et al., 2022), a new engine to replace Azure Durable Functions.
9. Conclusions
@@ Acknowledgments
Larissa Schmid is supported by the pilot program Core Informatics at KIT (KiKIT) of the Helmholtz Association (HGF). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 program (grant agreement PSAP, No. 101002047). We would also like to thank the Swiss National Supercomputing Centre (CSCS) for providing us with access to their HPC machine Ault. We thank Amazon Web Services for supporting this research with credits through the AWS Cloud Credit for Research, and Google Cloud Platform through the Google Cloud Research Credits program with the award GCP19980904.
@@
Appendix A Artifact Appendix
A.1. Abstract
Our artifact contains the implementation of SeBS-Flow, data, and analysis scripts. We provide the following components:
  • sebs-flow-implementation - Source code of the benchmark suite.
  • sebs-flow-artifact - Benchmarking results obtained for the paper together with Python plotting and analysis scripts used for data analysis.
A.2. Description & Requirements
A.2.1. How to access
A.2.2. Hardware dependencies
Unix system capable of sending HTTP requests to the cloud.
A.2.3. Software dependencies
  • Docker (at least 19)
  • Python 3.7+ with pip and venv
  • libcurl and its headers must be available to install pycurl
  • Standard Linux tools and zip installed
A.2.4. Benchmarks
We provide benchmarks and data used as part of the implementation in serverless-benchmarks/benchmarks/600.workflows and serverless-benchmarks/benchmarks-data/600.workflows.
A.3. Set-up
To install the benchmark suite with support for all platforms used for our evaluation, use ./install.py --aws --azure --gcp. This will create a virtual environment in python-venv, and install necessary Python dependencies and third-party dependencies. To use SeBS-Flow, the new Python virtual environment has to be activated: . python-venv/bin/activate. To deploy benchmarks to a platform, account credentials musst be supplied. See serverless-benchmarks/docs/platforms.md for details.
For measuring the execution of serverless workflows, we use a Redis instance deployed on a VM in the same cloud region as the workflow and its resources. See serverlessbenchmarks/docs/workflows.md for details.
A.4. Evaluation workflow
A.4.1. Major Claims
RQ1: Runtime. We evaluate the runtime differences between platforms by executing Experiment E1. Results are shown in Figure 8 and 9 and discussed in Section 7.2.
RQ2.1: Orchestration Overhead. We evaluate the overheads caused by cloud storage I/O via executing Experiment E3 (Figure 9), by parallel scheduling via executing Experiment E4 (Figure 11), and by the return payload via Experiment E5 (Figure 10(b)).
RQ2.2: Critical Path Discrepancy. We evaluate how the critical path is impacted by OS noise via executing Experiment E6 (Figure 15) and using data from E1 (Figure 14(b)14(c)) and by cold starts by analyzing the results from E1.
RQ3: Usability for Scientific Workflows. We evaluate how well serverless workflow orchestrations are suited for execution of scientific benchmarks by comparing execution times and scaling of the workflow on cloud platforms and an HPC system using data from E1 for 1000Genomes and with Experiments E7 and E8 (Figure 15).
RQ4: Pricing. We evaluate the differences in pricing between the cloud platforms by comparing execution cost of our application benchmarks using data from E1 (Figure 16).
A.4.2. Experiments
All configuration files used are provided per platform and benchmark executed as part of the artifact.
E1: Burst execution of application benchmarks. Execution of all application benchmarks. The paper uses 180 burst workflow executions per benchmark with 30 executions triggered at once.
E2: Warm execution of application benchmarks. Execution of Machine Learning and MapReduce benchmarks in warm mode. The paper collects 60 workflow executions with at least one warm function invocation per benchmark with 30 executions triggered at once.
E3: Parallel Download. Execution of the parallel download microbenchmark with 20 functions downloading a file in parallel, with filesizes from $2^{10}b$ to $2^{28}b$. The paper uses 30 burst executions triggered at once.
E4: Parallel Sleep. Execution of the parallel sleep microbenchmark with $2\leq N\leq 16$ functions and sleep durations of $1\leq T\leq 20$s. The paper uses 30 burst executions triggered at once.
E5: Function Chain. Execution of the function chain microbenchmark with 10 functions, with functions returning $2^{5}\leq M2^{18}b$ sent to the next function. The paper uses 30 warm executions.
E6: OS Noise. Execution of the selfish detour microbenchmark collecting $N=5000$ events, executed with memory configurations from 128MB to 2048MB. The paper uses 30 warm executions.
E7: Execution of 1000Genomes on HPC system. Execution of the 1000Genomes workflow on an HPC system. The paper uses five repetitions per configuration and a Intel(R) 6154@3.00GHz CPU. To reproduce the result that the execution on an HPC system is much faster, however, the workflow can be run in virtually every HPC environment.
E8: Scaling of individuals task of 1000Genomes workflow. Execution of the workflow 6101.1000-genome-individuals with inputs small-10 and small-20. The paper uses 180 burst workflow executions per input with 30 executions triggered at once.
A.5. Notes on Reusability
SeBS-Flow can be extended to evaluate workflow orchestrations on new serverless platforms. Moreover, new workflow benchmarks can be added to evaluate their performance on different platforms. Repetition of benchmarks over time can give insights into the evolution of performance.